OpenStudy (lovelyharmonics):

Find all polar coordinates of point P where P = ordered pair 5 comma pi divided by 3. @satellite73

OpenStudy (lovelyharmonics):

yay c: this ones really complicated!

OpenStudy (anonymous):

same as before locate \(\frac{\pi}{3}\) on the unit circle

OpenStudy (anonymous):

let me know when you see the coordinates, and we can finish in a second

OpenStudy (lovelyharmonics):

60degrees 1/2, sqrt{3}/2

OpenStudy (anonymous):

right

OpenStudy (lovelyharmonics):

the computer codes dont like me i guess .-.

OpenStudy (anonymous):

ignore the degrees, they have nothing to do with the problem now you have \(x=5\times \frac{1}{2}=\frac{5}{2}\)

OpenStudy (anonymous):

and \(y=5\times \frac{\sqrt3}{2}=\frac{5\sqrt3}{2}\)

OpenStudy (anonymous):

that is all

OpenStudy (anonymous):

what "computer codes"?

OpenStudy (lovelyharmonics):

.-. the sqrt{ } its suppose to make a squareroot sigh with out having to go under equation and everyting

OpenStudy (lovelyharmonics):

and if thats all then pick an answer cuz that isnt one of them.... (5, pi divided by 3 + nπ) or (-5, pi divided by 3 + nπ) (5, pi divided by 3 + (2n + 1)π) or (-5, pi divided by 3 + 2nπ) (5, pi divided by 3 + 2nπ) or (-5, pi divided by 3 + (2n + 1)π) (5, pi divided by 3 + 2nπ) or (-5, pi divided by 3 + 2nπ)

OpenStudy (anonymous):

ooh sorry, it said "find all" not "convert to rectangular" no problem

OpenStudy (lovelyharmonics):

lol thats why i said its so complicated in the beginning XD

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

let me look at the answers

OpenStudy (lovelyharmonics):

okay c:

OpenStudy (anonymous):

sine and cosine are periodic with period \(2\pi\) so one answer would be \[(5, \frac{\pi}{3}+2n\pi)\]

OpenStudy (anonymous):

then next part is a little hard for me to explain here but \(-5\) means go you off in the opposite direction, so you would have to add \(\pi\) to the angle to switch is back the other one will be \[(-5,\frac{\pi}{3}+(2n+1)\pi)\]

OpenStudy (anonymous):

that wasn't really english was it? well just go with C (it is always C)

OpenStudy (anonymous):

@ApollosChariot can you explain the last bit please

OpenStudy (anonymous):

@precal